OpenStudy (anonymous):

lcd of 3x/x^2-x-12Andx-4/x^2+3x

OpenStudy (anonymous):

\[3x divx ^{2}-x-12Andx-4divx ^{2}+3x\]

OpenStudy (mertsj):

\[\frac{3x}{(x-4)(x+3)} ,\frac{x-4}{x(x+3)}\]

OpenStudy (mertsj):

So we see that the factors of the denominators are x, x+3, x-4. All are to the first power. The LCD is the largest power of every different factor.

OpenStudy (anonymous):

Can u do one more?

OpenStudy (mertsj):

If it isn't too hard.

OpenStudy (anonymous):

:D ok

OpenStudy (anonymous):

What is the solution of x+3/x62-9+4/3-x-2=0

OpenStudy (anonymous):